# Schur–Weyl Equivalences for Wreath Product Superalgebras

### v1

@@ 2.5. Wreath Product Superalgebras
$(a_{1}\otimes b_{1})(a_{2}\otimes b_{2})=(-1)^{\overline{b_{1}}\cdot\overline{a_{1}}}(a_{1}a_{2})\otimes(b_{1}b_{2}).$
@@
In particular, $\mathds{k}S_{d}$ and $A^{\otimes d}$ are isomorphic to subsuperalgebras of $S_{d}\wr A$ in obvious ways. Also, note that there is a canonical surjective superalgebra map $S_{d}\wr A\to\mathds{k}S_{d}$.
@@
$\mathbf{v}_{\mathbf{i}}^{\mathbf{a}}\left(c_{1}\otimes\dotsb\otimes c_{d}\right)=(-1)^{\varepsilon(\mathbf{a},\mathbf{c})}\mathbf{v}_{\mathbf{i}}^{\mathbf{a}\cdot\mathbf{c}},$
for all $\mathbf{i}\in I(n,d)$, $\mathbf{a}\in I(A,d)$, and homogeneous $c_{1}\otimes\dotsb\otimes c_{d}\in A^{\otimes d}$, where $\mathbf{a}\cdot\mathbf{c}:=(a_{1}c_{1},\dotsc,a_{d}c_{d})$ and where $\varepsilon(\mathbf{a},\mathbf{c})=\sum_{1\leq p<q\leq d}\overline{a_{q}}\cdot\overline{c_{p}}$ is obtained by applying the rule of signs. Under this action $V_{n}^{\otimes d}$ is a free right $A^{\otimes d}$-supermodule with basis $\left\{\mathbf{v}^{\mathbf{1}}_{\mathbf{i}}\mid\mathbf{i}\in I(n,d)\right\}$. Also observe that this action can be extended to $S_{d}\wr A$ by having the elements of $S_{d}$ act by signed place permutation.
@@ 2.6. The Schur–Weyl Functor
Since $V_{n}^{\otimes d}$ is a $(U(\mathfrak{gl}_{n}(A),S_{d}\wr A)$-bisupermodule, there is a functor from left $S_{d}\wr A$-supermodules to left $U(\mathfrak{gl}_{n}(A))$-supermodules,
@@
As a special case, note that $F^{\mathds{k}}=F_{n,d}^{\mathds{k}}$ is the classical Schur–Weyl functor between $\mathds{k}S_{d}$-modules and $U(\mathfrak{gl}_{n}(\mathds{k}))$-modules. We will repeatedly use the well-known fact that $F_{n,d}^{\mathds{k}}$ gives an equivalence between the category of $\mathds{k}S_{d}$-supermodules and the category of polynomial supermodules of degree $d$ for $U(\mathfrak{gl}_{n}(\mathds{k}))$ whenever $n\geq d$, where both $\mathds{k}S_{d}$ and $U(\mathfrak{gl}_{n}(\mathds{k}))$ are viewed as superalgebras concentrated in parity $\bar{0}$ [18].
@@
$E^{\otimes d}\otimes_{\mathds{k}}N\to V^{\otimes d}\otimes_{A^{\otimes d}}N$ (2.6.2)
@@
Both superspaces are $\mathds{k}S_{d}$-supermodules where $S_{d}$ acts diagonally on each. The map given in (2.6.2) is an isomorphism of $\mathds{k}S_{d}$-supermodules and restricts to an isomorphism between the $S_{d}$-invariant subspaces. As a consequence,
$E^{\otimes d}\otimes_{\mathds{k}S_{d}}N\cong\left(E^{\otimes d}\otimes_{\mathds{k}}N\right)^{S_{d}}\cong\left(V^{\otimes d}\otimes_{A^{\otimes d}}N\right)^{S_{d}}\cong V^{\otimes d}\otimes_{S_{d}\wr A}N.$ (2.6.3)
Observe that these are $U(\mathfrak{gl}_{n}(\mathds{k}))$-supermodule isomorphisms. Going forward, we will identify these $U(\mathfrak{gl}_{n}(\mathds{k}))$-supermodules via this map. In particular, via this identification the square (2.6.1) commutes on objects. Likewise, it is straightforward to see that, up to applying (2.6.3), the square commutes on morphisms.
@@ 3. The Schur–Weyl Equivalence
is an equivalence of categories when $n\geq d+1$. As remarked after Definition 2.4.1, the image of $F^{A}_{n,d}$ lies in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$. Proposition 2.6.2 shows $F^{A}_{n,d}$ is faithful. Therefore, all that remains is to show that it is full and essentially surjective.
@@ 3.1. Key Injectivity Result
Let
$S=\mathds{k}-\text{span}\left\{v_{j_{1}}^{c_{1}}\otimes\dotsb\otimes v_{j_{d}}^{c_{d}}\mid\left\{j_{1},\dotsc,j_{d}\right\}=\left\{i_{1},\dotsc,i_{d}\right\},c_{1},\dotsc,c_{d}\in A\right\},$
and
$S^{\perp}=\mathds{k}-\text{span}\left\{v_{j_{1}}^{c_{1}}\otimes\dotsb\otimes v_{j_{d}}^{c_{d}}\mid\left\{j_{1},\dotsc,j_{d}\right\}\neq\left\{i_{1},\dotsc,i_{d}\right\},c_{1},\dotsc,c_{d}\in A\right\}.$
@@ 3.2. Fullness
Let $N_{1}$ and $N_{2}$ be $S_{d}\wr A$-supermodules and let $g\in\operatorname{Hom}_{U(\mathfrak{gl}_{n}(A))}\left(V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N_{1},V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N_{2}\right)$ be a homogeneous map. Applying the restriction functor and recalling the identification we make using (2.6.3), we may view $g$ as a morphism in $\operatorname{Hom}_{U(\mathfrak{gl}_{n}(\mathds{k}))}\left(E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N_{1},E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N_{2}\right)$. Since $F_{n,d}^{\mathds{k}}$ is a full functor, we may select a homogeneous $\mathds{k}S_{d}$-supermodule homomorphism $f:N_{1}\to N_{2}$ of the same parity as $g$ such that
@@
That is, $\operatorname{Id}_{V_{n}^{\otimes d}}\otimes f=g$ on the subspace of $V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N$ spanned by $\left\{v_{\mathbf{i}}^{\mathbf{1}}\otimes n\mid\mathbf{i}\in I(n,d),n\in N\right\}$. We claim that $f$ is actually a $S_{d}\wr A$-supermodule homomorphism. That is, $f$ is an allowable input into the functor $F_{n,d}^{A}$. This, along with the identification given in (2.6.3), demonstrates that $\mathcal{F}_{n,d}^{A}(f)=g$, as desired.
@@
By Theorem 3.1.1,
$f\left(c_{k}n\right)=(-1)^{\overline{g}\overline{E_{k,k}^{c}}}c_{k}f(n).$
@@ 3.3. Defining a $S_{d}\wr A$-module structure
Let $n\geq d+1$. Assume $N$ is a $\mathds{k}S_{d}$-supermodule for which $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ has a $U(\mathfrak{gl}_{n}(A))$-supermodule structure that makes it an object of $\mathcal{P}_{d}$ and, furthermore, that action restricts to the standard action of $U(\mathfrak{gl}_{n}(\mathds{k}))$ on $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$. Throughout this section, we assume we are in this situation. Given this setup it is possible to extend the $\mathds{k}S_{d}$-action on $N$ to all of $S_{d}\wr A$, as we now explain.
@@
$E_{j_{k},i_{k}}^{a}(\mathbf{v}_{\mathbf{i}}\otimes n)\in\left(E_{n}^{\otimes d}\otimes N\right)_{\boldsymbol{\varepsilon}_{\mathbf{j}}}=\left(E_{n}^{\otimes d}\right)_{\boldsymbol{\varepsilon}_{\mathbf{j}}}\otimes N.$
@@
$E_{j_{k},i_{k}}^{a}(\mathbf{v}_{\mathbf{i}}\otimes n)=\sum_{\sigma\in S_{d}}\mathbf{v}_{\mathbf{j}}\sigma\otimes n_{\sigma}=\mathbf{v}_{\mathbf{j}}\otimes\left(\sum_{\sigma\in S_{d}}\sigma n_{\sigma}\right)=\mathbf{v}_{\mathbf{j}}\otimes\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n).$
Since $\mathbf{j}$ has distinct entries, Theorem 3.1.1 and the above equality determines $\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n)$ and, hence, the function $\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}$. The fact that it is a $\mathds{k}$-linear map of parity $\overline{a}$ is straightforward to check.
The following result records some of the basic properties of the maps $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}:N\to N$. Note that here and going forward, we frequently make use of the observation that if $M$ is a supermodule for $U(\mathfrak{gl}_{n}(A))$ which lies in $\mathcal{P}_{d}$, and if $m\in M_{\lambda}$ and $\lambda(h_{j})=0$, then $E_{i,j}^{a}m=0$: either because $i=j$ and the second condition of Definition 2.4.1 applies, or because $i\neq j$ and because $\lambda+\varepsilon_{i}-\varepsilon_{j}$ is not in $\Lambda^{+}(n,d)$ and, hence, is not an allowable weight of $M$.
@@
    If $(\mathbf{j},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$, then for any $p\in[1,n]\setminus\{i_{1},\dotsc,i_{d},j_{k}\}$ and $\ell\in[1,d]$ with $\ell\neq k$, if we define $\mathbf{i}^{\prime}=(i_{1},\dotsc,i_{\ell-1},p,i_{\ell+1},\dotsc,i_{d})$ and $\mathbf{j}^{\prime}=(j_{1},\dotsc,j_{\ell-1},p,j_{\ell+1},\dotsc,j_{d})$, then $(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$ and $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}}$ for all homogeneous $a\in A$.
@@
We first consider the special case when it is the pair $(\mathbf{j},\mathbf{i})_{k}$ and $(\mathbf{i},\mathbf{i})_{k}$. The statement is trivial if $\mathbf{j}=\mathbf{i}$, so we assume $\mathbf{j}\neq\mathbf{i}$ and, hence, $j_{k}\neq i_{k}$. Use the supercommutator formula to see that
@@
However, $\boldsymbol{\varepsilon}_{\mathbf{i}}+\varepsilon_{i_{k}}-\varepsilon_{j_{k}}\not\in\Lambda^{+}(n,d)$ and so the first term of the final sum equals zero; likewise, the observation that $h_{j_{k}}\mathbf{v}_{\mathbf{i}}=0$ along with the second property of supermodules in $\mathcal{P}_{d}$ implies the last term is also equal to zero. Therefore, $\mathbf{v}_{\mathbf{i}}\otimes\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}(n)=\mathbf{v}_{\mathbf{i}}\otimes\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}(n)$. It follows from Theorem 3.1.1 that $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}(n)=\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}(n)$, as desired.
The general case follows by applying the special case to the pairs $(\mathbf{j},\mathbf{i})_{k}$ and $(\mathbf{i},\mathbf{i})_{k}$, and $(\mathbf{j}^{\prime},\mathbf{i})_{k}$ and $(\mathbf{i},\mathbf{i})_{k}$ to obtain $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{j}^{\prime},\mathbf{i})}$.
Likewise, to prove (2), it suffices to consider the case of $(\mathbf{j},\mathbf{i})_{k}$ and $(\mathbf{j},\mathbf{j})_{k}$ where $\mathbf{j}\neq\mathbf{i}$. In particular, $j_{k}\neq i_{k}$. Use the supercommutator formula to see that
@@
However, the first term in the last sum is again zero by the second property of supermodules in $\mathcal{P}_{d}$. Theorem 3.1.1 then implies $\alpha_{(\mathbf{j},\mathbf{j})_{k}}^{a}(n)=\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n)$, as claimed.
To prove (3), we first note that, by the choice of $p$, the entries of $\mathbf{i}^{\prime}$ and $\mathbf{j}^{\prime}$ are pairwise distinct and the two have the same values in every position except possibly at $i^{\prime}_{k}$ and $j^{\prime}_{k}$. Therefore $(\mathbf{i}^{\prime},\mathbf{j}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$, as claimed. Noting that $j^{\prime}_{k}=j_{k}$ and $i^{\prime}_{k}=i_{k}$, and, hence, that $i^{\prime}_{k}\neq p$ and $j^{\prime}_{k}\neq i_{\ell}$, use the supercommutator formula to see that
@@
Theorem 3.1.1 implies $\alpha_{(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}}^{a}(n)=\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n)$, as claimed. ∎
@@
Let $n\geq d$. For $(\mathbf{j},\mathbf{i})_{k},(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$ and homogeneous $a\in A$,
@@
Since the statement is trivially true if $n=d$, we assume $n\geq d+1$ for the remainder of the proof.

### v2

@@ 2.5. Wreath Product Superalgebras
$(a_{1}\otimes b_{1})(a_{2}\otimes b_{2})=(-1)^{\overline{b_{1}}\cdot\overline{a_{2}}}(a_{1}a_{2})\otimes(b_{1}b_{2}).$
@@
In particular, $\mathds{k}S_{d}$ and $A^{\otimes d}$ are isomorphic to subsuperalgebras of $S_{d}\wr A$ in obvious ways.
@@
$\mathbf{v}_{\mathbf{i}}^{\mathbf{a}}\left(c_{1}\otimes\dotsb\otimes c_{d}\right)=(-1)^{\gamma(\mathbf{a},\mathbf{c})}\mathbf{v}_{\mathbf{i}}^{\mathbf{a}\cdot\mathbf{c}},$
for all $\mathbf{i}\in I(n,d)$, $\mathbf{a}\in I(A,d)$, and homogeneous $c_{1}\otimes\dotsb\otimes c_{d}\in A^{\otimes d}$, where $\mathbf{a}\cdot\mathbf{c}:=(a_{1}c_{1},\dotsc,a_{d}c_{d})$ and where $\gamma(\mathbf{a},\mathbf{c})=\sum_{1\leq p<q\leq d}\overline{a_{q}}\cdot\overline{c_{p}}$ is obtained by applying the rule of signs. Under this action $V_{n}^{\otimes d}$ is a free right $A^{\otimes d}$-supermodule with basis $\left\{\mathbf{v}^{\mathbf{1}}_{\mathbf{i}}\mid\mathbf{i}\in I(n,d)\right\}$. Also observe that this action can be extended to $S_{d}\wr A$ by having the elements of $S_{d}$ act by signed place permutation. It is straightforward to verify the actions of $U(\mathfrak{gl}_{n}(A))$ and $S_{d}\wr A$ on $V_{n}^{\otimes d}$ mutually commute.
@@ 2.6. The Schur–Weyl Functor
Since $V_{n}^{\otimes d}$ is a $(U(\mathfrak{gl}_{n}(A)),S_{d}\wr A)$-bisupermodule, there is a functor from left $S_{d}\wr A$-supermodules to left $U(\mathfrak{gl}_{n}(A))$-supermodules,
@@
As a special case, note that $F^{\mathds{k}}=F_{n,d}^{\mathds{k}}$ is the classical Schur–Weyl functor between $\mathds{k}S_{d}$-modules and $U(\mathfrak{gl}_{n}(\mathds{k}))$-modules. We will repeatedly use the well-known fact that $F_{n,d}^{\mathds{k}}$ gives an equivalence between the category of $\mathds{k}S_{d}$-supermodules and the category of polynomial supermodules of degree $d$ for $U(\mathfrak{gl}_{n}(\mathds{k}))$ whenever $n\geq d$, where both $\mathds{k}S_{d}$ and $U(\mathfrak{gl}_{n}(\mathds{k}))$ are viewed as superalgebras concentrated in parity $\bar{0}$ [Green, 2007, Chapter 6].
@@
$E_{n}^{\otimes d}\otimes_{\mathds{k}}N\to V_{n}^{\otimes d}\otimes_{A^{\otimes d}}N$ (2.6.2)
@@
Both $E_{n}^{\otimes d}\otimes_{\mathds{k}}N$ and $V_{n}^{\otimes d}\otimes_{A^{\otimes d}}N$ are $\mathds{k}S_{d}$-supermodules where $S_{d}$ acts diagonally on each. The map given in (2.6.2) is $S_{d}$-equivariant. Hence, it is an isomorphism of $\mathds{k}S_{d}$-supermodules and restricts to an isomorphism between the $S_{d}$-invariant subspaces. With this observation we have the middle isomorphism in the following chain:
$E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N\cong\left(E_{n}^{\otimes d}\otimes_{\mathds{k}}N\right)^{S_{d}}\cong\left(V_{n}^{\otimes d}\otimes_{A^{\otimes d}}N\right)^{S_{d}}\cong V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N.$ (2.6.3)
The first and last isomorphisms are straightforward. Observe that all three maps are $U(\mathfrak{gl}_{n}(\mathds{k}))$-supermodule isomorphisms. Going forward, we will identify $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ and $V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N$ as $U(\mathfrak{gl}_{n}(\mathds{k}))$-supermodules via the composite of these maps.
In particular, using this identification one sees that the square (2.6.1) commutes for objects: $res_{U(\mathfrak{gl}_{n}(\mathds{k}))}^{U(\mathfrak{gl}_{n}(A))}\left(F_{n,d}^{A}(N)\right)=res_{\mathds{k}S_{d}}^{S_{d}\wr A}\left(F_{n,d}^{\mathds{k}}(N)\right)$ for all $S_{d}\wr A$-supermodules $N$. Likewise, it is straightforward to see that the square commutes on morphisms.
@@ 3. The Schur–Weyl Equivalence
is an equivalence of categories when $n\geq d+1$. As remarked after Definition 2.4.1, the image of $F^{A}_{n,d}$ lies in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$. Proposition 2.6.2 shows $F^{A}_{n,d}$ is faithful. Therefore, all that remains is to show it is full and essentially surjective.
@@ 3.1. Key Injectivity Result
Let $S$ be the $\mathds{k}$-span of the set
$\left\{v_{j_{1}}^{c_{1}}\otimes\dotsb\otimes v_{j_{d}}^{c_{d}}\mid\left\{j_{1},\dotsc,j_{d}\right\}=\left\{i_{1},\dotsc,i_{d}\right\},c_{1},\dotsc,c_{d}\in A\right\},$
and let $S^{\perp}$ be the $\mathds{k}$-span of the set
$\left\{v_{j_{1}}^{c_{1}}\otimes\dotsb\otimes v_{j_{d}}^{c_{d}}\mid\left\{j_{1},\dotsc,j_{d}\right\}\neq\left\{i_{1},\dotsc,i_{d}\right\},c_{1},\dotsc,c_{d}\in A\right\}.$
@@ 3.2. Fullness
Let $N_{1}$ and $N_{2}$ be $S_{d}\wr A$-supermodules and let $g\in\operatorname{Hom}_{U(\mathfrak{gl}_{n}(A))}\left(V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N_{1},V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N_{2}\right)$ be a homogeneous map. Applying the restriction functor and the identification given by (2.6.3), we may view $g$ as a morphism in $\operatorname{Hom}_{U(\mathfrak{gl}_{n}(\mathds{k}))}\left(E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N_{1},E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N_{2}\right)$. Since $F_{n,d}^{\mathds{k}}$ is a full functor, we may select a homogeneous $\mathds{k}S_{d}$-supermodule homomorphism $f:N_{1}\to N_{2}$ of the same parity as $g$ such that
@@
That is, $\operatorname{Id}_{V_{n}^{\otimes d}}\otimes f=g$ on the subspace of $V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N$ spanned by $\left\{v_{\mathbf{i}}^{\mathbf{1}}\otimes n\mid\mathbf{i}\in I(n,d),n\in N\right\}$. We claim that $f$ is actually a $S_{d}\wr A$-supermodule homomorphism. That is, $f$ is an allowable input into the functor $F_{n,d}^{A}$. This, along with the identification given in (2.6.3), demonstrates that $F_{n,d}^{A}(f)=g$, as desired.
@@
By Proposition 3.1.1,
$f\left(c_{k}n\right)=(-1)^{\overline{g}\cdot\overline{E_{k,k}^{c}}}c_{k}f(n).$
@@ 3.3. Defining a $S_{d}\wr A$-module structure
Let $n\geq d+1$. Assume $N$ is a $\mathds{k}S_{d}$-supermodule for which $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ has a $U(\mathfrak{gl}_{n}(A))$-supermodule structure that makes it an object of $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$ and, furthermore, that action restricts to the standard action of $U(\mathfrak{gl}_{n}(\mathds{k}))$ on $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$. Throughout this section, we assume we are in this situation.
The goal of this section is to show that with this setup it is possible to extend the $\mathds{k}S_{d}$-action on $N$ to an action by $S_{d}\wr A$ on $N$. The general strategy is to use the assumed action of $\mathfrak{gl}_{n}(A)$ on $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ to define the action of $A^{\otimes d}$ on $N$. Very roughly, given $a\in A$ and $n\in N$, we would like to define the action of $a_{k}:=1^{\otimes(k-1)}\otimes a\otimes 1^{\otimes d-k}$ on $n$ using the action of $E_{j_{k},i_{k}}^{a}$ on $v_{\mathbf{i}}\otimes n$ for $(\mathbf{j},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$ in combination with Proposition 3.1.1. Let us explain how to implement this strategy.
@@
$E_{j_{k},i_{k}}^{a}(\mathbf{v}_{\mathbf{i}}\otimes n)\in\left(E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N\right)_{\boldsymbol{\varepsilon}_{\mathbf{j}}}=\left(E_{n}^{\otimes d}\right)_{\boldsymbol{\varepsilon}_{\mathbf{j}}}\otimes_{\mathds{k}S_{d}}N.$
@@
$E_{j_{k},i_{k}}^{a}(\mathbf{v}_{\mathbf{i}}\otimes n)=\sum_{\sigma\in S_{d}}\left(\mathbf{v}_{\mathbf{j}}\sigma\otimes n_{\sigma}\right)=\mathbf{v}_{\mathbf{j}}\otimes\left(\sum_{\sigma\in S_{d}}\sigma n_{\sigma}\right)=\mathbf{v}_{\mathbf{j}}\otimes\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n),$
for some $\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n)\in N$. Since $\mathbf{j}$ has distinct entries, Proposition 3.1.1 implies $\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n)$ is unique and, hence, the rule $n\mapsto\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n)$ defines a function, as claimed. The fact that it is a $\mathds{k}$-linear map of parity $\overline{a}$ is straightforward to check.
The following result records some of the basic properties of the maps $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}:N\to N$. Note that here and going forward, we frequently make use of the observation that if $M$ is a supermodule for $U(\mathfrak{gl}_{n}(A))$ which lies in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$, and if $m\in M_{\lambda}$ and $\lambda(h_{j})=0$, then $E_{i,j}^{a}m=0$: either because $i=j$ and the second condition of Definition 2.4.1 applies, or because $i\neq j$ and because $\lambda+\varepsilon_{i}-\varepsilon_{j}$ is not in $\Lambda(n,d)$ and, hence, is not an allowable weight of $M$.
@@
    If $(\mathbf{j},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$, then for any $p\in[1,n]\setminus\{i_{1},\dotsc,i_{d},j_{1},\dotsc,j_{d}\}$ and $\ell\in[1,d]$ with $\ell\neq k$, if we define $\mathbf{i}^{\prime}=(i_{1},\dotsc,i_{\ell-1},p,i_{\ell+1},\dotsc,i_{d})$ and $\mathbf{j}^{\prime}=(j_{1},\dotsc,j_{\ell-1},p,j_{\ell+1},\dotsc,j_{d})$, then $(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$ and $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}}$ for all homogeneous $a\in A$.
@@
To prove (1), we first consider the special case of the pair $(\mathbf{j},\mathbf{i})_{k}$ and $(\mathbf{i},\mathbf{i})_{k}$. The claim in (1) is trivial in this case if $\mathbf{j}=\mathbf{i}$, so we assume $\mathbf{j}\neq\mathbf{i}$ and, hence, $j_{k}\neq i_{k}$. Use the supercommutator formula (2.2.1) to see that
@@
However, $\boldsymbol{\varepsilon}_{\mathbf{i}}+\varepsilon_{i_{k}}-\varepsilon_{j_{k}}\not\in\Lambda(n,d)$ and so the first term of the final sum equals zero; likewise, the observation that $h_{j_{k}}\mathbf{v}_{\mathbf{i}}=0$ along with the second property of supermodules in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$ implies the last term is also equal to zero. Therefore, $\mathbf{v}_{\mathbf{i}}\otimes\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}(n)=\mathbf{v}_{\mathbf{i}}\otimes\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}(n)$. It follows from Proposition 3.1.1 that $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}(n)=\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}(n)$, as desired.
Applying the special case to the pair $(\mathbf{j},\mathbf{i})_{k}$ and $(\mathbf{i},\mathbf{i})_{k}$, and the pair $(\mathbf{j}^{\prime},\mathbf{i})_{k}$ and $(\mathbf{i},\mathbf{i})_{k}$, yields $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}$ and $\alpha^{a}_{(\mathbf{j}^{\prime},\mathbf{i})}=\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}$, respectively. Combining the two gives $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{j}^{\prime},\mathbf{i})}$, as desired.
Likewise, to prove (2), it suffices to consider the case of $(\mathbf{j},\mathbf{i})_{k}$ and $(\mathbf{j},\mathbf{j})_{k}$ where $\mathbf{j}\neq\mathbf{i}$. In particular, $j_{k}\neq i_{k}$. Use the supercommutator formula (2.2.1) to see that
@@
However, the first term in the last sum is again zero by the second property of supermodules in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$. Proposition 3.1.1 then implies $\alpha_{(\mathbf{j},\mathbf{j})_{k}}^{a}(n)=\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n)$, as claimed.
To prove (3), we first note that, by the choice of $p$, the entries of $\mathbf{i}^{\prime}$ and $\mathbf{j}^{\prime}$ are pairwise distinct and the two have the same values in every position except possibly at $i^{\prime}_{k}$ and $j^{\prime}_{k}$. Therefore $(\mathbf{i}^{\prime},\mathbf{j}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$, as claimed. Noting that $j^{\prime}_{k}=j_{k}$ and $i^{\prime}_{k}=i_{k}$, and, hence, that $i^{\prime}_{k}\neq p$ and $j^{\prime}_{k}\neq i_{\ell}$, the supercommutator formula (2.2.1) implies that
@@
Proposition 3.1.1 implies $\alpha_{(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}}^{a}(n)=\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(n)$, as claimed. ∎
@@
Let $n\geq d+1$. For $(\mathbf{j},\mathbf{i})_{k},(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$ and homogeneous $a\in A$,